# CAST-Phys: Contactless Affective States Through Physiological signals Database

### v1

@@ IV-A2 Facial Expressions
From the facial videos recorded during the experiment sessions, we extract facial expression features using the Py-Feat estimator [76]. Given the high frame rate of the recordings, we sample facial features at one-second intervals throughout each trial video. For each second of the video, we extract 20 facial features based on the Facial Action Coding System (FACS) [77], one of the most widely used frameworks for systematically quantifying facial muscle movements. FACS identifies the actions of distinct groups of facial muscles, referred to as action units (AUs), which are recognized as reliable indicators of emotional expression [78].
@@ IV-B Self-report annotations
Figure 6 presents the mean self-report results from the CAST-Phys dataset. For each of the nine elicited quadrants, two dots represent the two videos used to stimulate each quadrant. Analyzing the results from 61 participants, we find that the elicitation strategy was effective, leading to a diverse emotional representation. Overall, there is strong agreement across most quadrants; however, variations are observed, particularly concerning arousal. While valence closely aligns with the desired elicitation, some discrepancies in arousal are noted. For example, quadrant seven, representing low arousal and low valence, shows a higher arousal rating similar to quadrant four. Conversely, quadrant five yielded slightly lower arousal ratings from both videos. For the valence dimension, the average responses are generally accurate, except for one video from quadrant three, which produced a more neutral response. Nonetheless, averaging the self-report data can hide key behaviors essential for understanding the emotional response in our dataset. Consequently, we conduct a more in-depth analysis, as shown in Figures 7 and 8.
Figure 7 displays the self-assessment frequencies (number of times that the quadrant was annotated) for the nine elicited quadrants, as well as separate frequencies for valence and arousal. The nine-quadrant distribution reveals an imbalance in the annotations: specifically, quadrant 2 (representing high arousal with neutral valence) is the least populated one, while quadrants 1 and 3 account for the majority of annotations. This suggests that although quadrant 2 successfully elicited the intended arousal response, the valence outcome was split between positive and negative, yielding an average neutral effect as seen in Figure 6. Additionally, quadrant 8 shows the third highest frequency, consistent with the observation that the two videos originally intended for quadrant five were instead annotated as quadrant 8. The rest of the observed quadrants show a very similar distribution. Following [12], we computed the Cohen’s Kappa coefficient for the nine categorical annotations, obtaining an average of 0.3463, similar to [12]. We further analyzed the contributions of valence and arousal separately by categorizing each into three classes: positive, neutral, and negative for valence, and activated, neutral, and deactivated for arousal. In both cases, the neutral state is the least represented one, even though the overall distribution across classes is more balanced than using nine classes, with positive and activated responses being the most common.
@@
These results suggest that some emotions are inherently more difficult to elicit than others, as individuals often respond differently to the same stimuli, even assigning opposite valence or arousal ratings. Factors such as education and cultural background may contribute to the variability in annotation agreement. Moreover, the intensity of emotional responses can differ among users. Nowadays, our society is exposed to a large amount of multimedia content, and many individuals are accustomed to high levels of stimulation, meaning that some may require stronger stimuli to experience emotions more intensely.
@@ IV-C Multi-modal emotion recognition
After applying the ANOVA test, we identify 26 meaningful features for arousal and 15 for valence:
  • For the arousal dimension, the PPG modality yields two temporal features (heart rate standard deviation and SDNN), two frequency features (LF and the LF/HF ratio), and three non-linear features (TINN, SD2, and the Poincaré ratio between SD1 and SD2). The respiration signal produces significant results for all features except the mean of respiration derivation. In the EDA modality, only the SCL mean and SCR mean of peak magnitudes are not significant. Facial expressions provide 11 meaningful features, including AU02, AU04, AU06, AU10, AU12, AU14, AU15, AU17, AU25, AU26, and AU28.
  • For the valence dimension, the PPG modality retains only two significant features, LF and the LF/HF ratio. The EDA modality contributes four significant features: SCL and SCR standard deviations, SCR mean, and the mean of peak magnitudes. Facial expressions are the most significant modality, including AU01, AU06, AU09, AU12, AU14, AU17, AU24, AU25, and AU26. The respiration signal does not yield any significant features for valence. However, the respiration mean is considered only for single-modality computation, as its p-value of 0.067 is close to 0.05, approaching the margin of significance.
@@
Figure 9 presents the emotion recognition results, showing the best accuracy for each modality, as well as their fusion for both arousal (on the left) and valence (on the right). Regarding single-modality performance, for arousal dimension, heart rate and respiration modalities demonstrate the highest performance, with 50.1% accuracy using AdaBoost and 49.3% accuracy using Random Forest, respectively. Facial expression features achieve 45% accuracy with SVM, while EDA features reach 41% accuracy with Random Forest. For the valence dimension, the heart rate features reach 49.6% accuracy with Adaboost, followed by respiration and EDA, with 47.7 % and 47.2%, respectively. Finally, faces obtain a performance of 42.9 % with Naive Bayes.
Surprisingly, facial expressions yielded the lowest valence classification performance. This result is unexpected, given that facial expressions are strong indicators of valence and previously exhibited the highest number of meaningful features. To investigate potential errors in the AUs model, we conducted a manual annotation experiment where three external annotators independently labelled the same frames used for the facial model. The annotators achieved a valence accuracy of 45%, closely matching the model’s performance.
@@
Figure 10: Facial expressions of four subjects during stimulus $Q3\_2$ at the same moment. All subjects reported positive and activated emotional responses ($Q3$).
Finally, the fusion of modalities (represented in magenta) significantly enhances performance compared to single-modality approaches for both arousal and valence. By integrating all four recorded modalities, we achieve an accuracy of 52.6% for arousal and 54.7% for valence, representing an improvement of 2.5% and 5%, respectively, over the best-performing single modality. For a more detailed analysis, Table II presents the six-fold cross-validation results for emotion classification across different classifiers within the fusion-based model. The findings indicate that Random Forest achieves the highest accuracy and F1-score for both arousal and valence, followed closely by AdaBoost and Decision Tree classifiers.
@@
[…]
@@ IV-D Remote physiological signal extraction
Considering the results and importance of heart rate and respiration modalities in the emotion recognition of the previous subsection, we decide to extract these two signals remotely due to their potential link with emotional cues. To extract physiological signals from facial videos, we compare three state-of-the-art methods: two hand-crafted approaches, ICA [29] and POS [32], and one data-driven method, TDM [47]. These methods are chosen for their versatility, efficiency and suitability for real-time applications while maintaining competitive heart rate (HR) estimation performance. Table III presents the six-fold average performance of each method using common remote HR and respiration rate (RR) estimation metrics, including mean absolute error (MAE), root mean squared error (RMSE), and Pearson correlation ($\rho$). The results indicate that the hand-crafted models yield acceptable performance, with the POS method achieving an error of 2.01 BPM across the entire dataset. However, the data-driven TDM approach outperforms both, obtaining the lowest HR error with an MAE of 1.1 BPM, an RMSE of 4.51 BPM, and a correlation of 0.93. Additionally, although the original TDM model was designed exclusively for rPPG signal recovery, we enhanced it by, incorporating two final fully connected layers, yielding a multi-task regression framework that simultaneously estimates both rPPG and rRSP signals. The model was trained on the CAST-Phys dataset following the same six-fold cross-validation strategy described in Subsection IV-C. For remote respiration estimation, we achieved baseline results of 2.21 RPM MAE and 3.25 RPM RMSE. Given its superior performance and ability to extract both PPG and respiration signals, we employ the TDM model to obtain the remote physiological signals used for emotion recognition, as detailed in the next subsection. Figure 11 depicts an example of our remote physiological extraction during the high-arousal elicitation observed at the end of the video.
@@ IV-E Remote multi-modal emotion recognition
The selected features closely resemble those identified in Subsection IV-C. For the arousal dimension, heart rate yields four significant features: pNN50, heart rate standard deviation, LF, and the LF/HF ratio, while respiration retains the same three meaningful features as before. Regarding valence, heart rate includes four relevant features: LF, LF/HF ratio, TINN, and the Poincaré ratio between SD1 and SD2. Additionally, respiration now contributes two meaningful features: respiration rate mean and standard deviation.
@@
Figure 12 presents the accuracy performance of remote emotion recognition for each modality. For arousal, heart rate exhibits a notable performance drop, achieving only 40.7% accuracy with Random Forest compared to the contact-based signal. In contrast, respiration maintains a similar performance, reaching 47.6% accuracy with Random Forest, closely matching the contact-based results. For valence, the performance remains nearly unchanged between remote and contact-based modalities. Heart rate achieves 51.1% accuracy using Random Forest, slightly outperforming the 49.6% recorded in the contact-based setup. Similarly, respiration attains 46.3% accuracy with Random Forest, compared to 47.6% in the contact-based condition. These findings suggest that remote emotion recognition can achieve comparable results to contact-based methods, with the primary discrepancy observed in arousal estimation using remote heart rate features.
@@
TABLE IV: Remote multi-modal emotion recognition 6-Fold Average Performance
[…]
@@
The drop in arousal accuracy for remote heart rate can be attributed to the fewer features extracted by ANOVA, particularly the absence of important non-linear features that contribute to higher accuracy. In contrast, valence performance benefits from the same two frequency-domain features used in the contact-based approach, along with two additional non-linear features, leading to a 1.5% improvement in accuracy. Hence, these results also highlight the critical role of feature selection in emotion recognition, as heart rate variability non-linear features appear to significantly impact both arousal and valence estimation. Exploring alternative feature selection methods or employing data-driven feature selection techniques could further enhance these baseline results.
@@
As shown in Subsection IV-C, the combination of the extracted modalities, in this case with facial expressions, rPPG and rRSP, reflects in Figure 12 and Table IV a superior performance than each modality individually. From Table IV, we appreciate that Random Forest, Decision Tree and Adaboost seem to be the most robust classifiers for the evaluated modalities like in contact-based estimation. Comparing the multi-modal emotion recognition results obtained from contact and fully remote modalities, can observe very similar performance with a slight decrease in remote modality. Figure 13 presents the confusion matrices for valence and arousal classification across different cases. For valence (Figures 13c and 13d), we observe a common issue in both contact-based and remote approaches: the neutral state is not well detected, and valence tends to be polarized between positive and negative emotions. This aligns with the findings discussed in previous subsections, where the presence of neutral states within positive and negative emotions makes accurate classification of this category more challenging. In contrast, for arousal (Figures 13a and 13b), the neutral class achieves significantly better classification in both cases, with even stronger performance in remote emotion recognition. Although remote emotion recognition shows slightly lower accuracy for activated and deactivated states compared to contact-based methods, it appears to achieve a more balanced classification across the three arousal categories.
(a) Arousal multi-modal contact-based
(b) Arousal multi-modal remote-based
(c) Valence multi-modal contact-based
(d) Valence multi-modal remote-based
@@ V Discussion and future work
While this study presents the first approach to remote emotion estimation using three different modalities (rPPG, rRSP, facial expressions), EDA was not incorporated, as it remains largely unexplored in remote sensing. Recent studies [79, 80] have demonstrated the feasibility of recovering EDA signals through camera-based sensing. The integration of remote EDA into multi-modal emotion recognition frameworks could significantly enhance the estimation of sympathetic responses alongside modalities like rPPG or rRSP. In this sense, CAST-Phys can be a valuable dataset for such research lines, both in terms of providing training data and serving as a public benchmark.

### v2

@@ IV-A2 Facial Expressions
From the facial videos recorded during the experiment sessions, we extract facial expression features using the Py-Feat estimator [cheong2023py]. Given the high frame rate of the recordings, we sample facial features at one-second intervals throughout each trial video. For each second of the video, we extract 20 facial features based on the Facial Action Coding System (FACS) [ekman1978facial], one of the most widely used frameworks for systematically quantifying facial muscle movements. FACS identifies the actions of distinct groups of facial muscles, referred to as action units (AUs), which are recognized as reliable indicators of emotional expression [sayette2001psychometric].
@@
(a) Valence mean $\pm$ SD per intended quadrant.
(b) Arousal mean $\pm$ SD per intended quadrant.
Figure 8: Mean and standard deviation of (a) valence and (b) arousal ratings across the nine intended quadrants.
@@ IV-B Self-report annotations
Figure 6 presents the mean self-report results from the CAST-Phys dataset. For each of the nine elicited quadrants, two dots represent the two videos used to stimulate each quadrant. Analyzing the results from 61 participants, we find that the elicitation strategy was effective, leading to a diverse emotional representation. Overall, there is strong agreement across most quadrants; however, variations are observed, particularly concerning arousal. While valence closely aligns with the desired elicitation, some discrepancies in arousal are noted. For example, quadrant seven, representing low arousal and low valence, shows a higher arousal rating similar to quadrant four. Conversely, quadrant five yielded slightly lower arousal ratings from both videos. For the valence dimension, the average responses are generally accurate, except for one video from quadrant three, which produced a more neutral response. Nonetheless, averaging the self-report data can hide key behaviors essential for understanding the emotional response in our dataset. Consequently, we conduct a more in-depth analysis, as shown in Figures 7, 8 and 9.
Figure 7 displays the self-assessment frequencies (number of times that the quadrant was annotated) for the nine elicited quadrants, as well as separate frequencies for valence and arousal. The nine-quadrant distribution reveals an imbalance in the annotations: specifically, quadrant 2 (representing high arousal with neutral valence) is the least populated one, while quadrants 1 and 3 account for the majority of annotations. This suggests that although quadrant 2 successfully elicited the intended arousal response, the valence outcome was split between positive and negative, yielding an average neutral effect as seen in Figure 6. Additionally, quadrant 8 shows the third highest frequency, consistent with the observation that the two videos originally intended for quadrant five were instead annotated as quadrant 8. The rest of the observed quadrants show a very similar distribution. Following [soleymani2011multimodal], we computed the Cohen’s Kappa coefficient for the nine categorical annotations, obtaining an average of 0.3463, similar to [soleymani2011multimodal]. We further analyzed the contributions of valence and arousal separately by categorizing each into three classes: positive, neutral, and negative for valence, and activated, neutral, and deactivated for arousal. In both cases, the neutral state is the least represented one, even though the overall distribution across classes is more balanced than using nine classes, with negative and activated responses being the most common.
To statistically validate the effectiveness of the emotion elicitation protocol, a one-way ANOVA was conducted on participants’ valence and arousal self-reports across the nine intended quadrants. The analysis revealed significant differences for both valence ($F(8,1089)=250.6,p<0.001$) and arousal ($F(8,1089)=110.4,p<0.001$) across the nine intended quadrants, confirming that the selected stimuli successfully elicited distinct affective responses. Figures 8a and 8b present the mean ± SD valence and arousal ratings per quadrant, highlighting the clear differentiation of emotional responses across the affective circumplex.
@@
These results suggest that some emotions are inherently harder to elicit, as individuals may react differently to the same stimuli, even assigning opposite valence or arousal ratings. Factors such as education and cultural background may contribute to the variability in annotation agreement. Moreover, the intensity of emotional responses can differ among users. Nowadays, our society is exposed to a large amount of multimedia content, and many individuals are accustomed to high levels of stimulation, meaning that some may require stronger stimuli to experience emotions more intensely.
@@ IV-C Multi-modal emotion recognition
After applying the ANOVA test, we identify 26 meaningful features for valence and 15 for arousal:
  • For the valence dimension, the PPG modality yields two temporal features (heart rate standard deviation and SDNN), two frequency features (LF and the LF/HF ratio), and three non-linear features (TINN, SD2, and the Poincaré ratio between SD1 and SD2). The respiration signal produces significant results for all features except the mean of respiration derivation. In the EDA modality, only the SCL mean and SCR mean of peak magnitudes are not significant. Facial expressions provide 11 meaningful features, including AU02, AU04, AU06, AU10, AU12, AU14, AU15, AU17, AU25, AU26, and AU28.
  • For the arousal dimension, the PPG modality retains only two significant features, LF and the LF/HF ratio. The EDA modality contributes four significant features: SCL and SCR standard deviations, SCR mean, and the mean of peak magnitudes. Facial expressions are the most significant modality, including AU01, AU06, AU09, AU12, AU14, AU17, AU24, AU25, and AU26. The respiration signal does not yield any significant features for valence. However, the respiration mean is considered only for single-modality computation, as its p-value of 0.067 is close to 0.05, approaching the margin of significance.
@@
Figure 10 presents the emotion recognition results, showing the best accuracy for each modality, as well as their fusion for both arousal (on the left) and valence (on the right). Regarding single-modality performance, for valence dimension, heart rate and respiration modalities demonstrate the highest performance, with 50.1% accuracy using AdaBoost and 49.3% accuracy using Random Forest, respectively. Facial expression features achieve 45% accuracy with SVM, while EDA features reach 41% accuracy with Random Forest. For the arousal dimension, the heart rate features reach 49.6% accuracy with Adaboost, followed by respiration and EDA, with 47.7 % and 47.2%, respectively. Finally, faces obtain a performance of 42.9 % with Naive Bayes.
Surprisingly, facial expressions yielded a low valence classification performance. This result is unexpected, given that facial expressions are strong indicators of valence and previously exhibited the highest number of meaningful features. To investigate potential errors in the AUs model, we conducted a manual annotation experiment where three external annotators independently labelled the same frames used for the facial model. The annotators achieved a valence accuracy of 45%, closely matching the model’s performance.
@@
Figure 11: Facial expressions of four subjects during stimulus Q3_2Q3\_2 at the same moment. All subjects reported positive and activated emotional responses ($Q3$).
Finally, the fusion of modalities (represented in magenta) significantly enhances performance compared to single-modality approaches for both valence and arousal. By integrating all four recorded modalities, we achieve an accuracy of 52.6% for valence and 54.7% for arousal, representing an improvement of 2.5% and 5%, respectively, over the best-performing single modality. For a more detailed analysis, Table II presents the six-fold cross-validation results for emotion classification across different classifiers within the fusion-based model. The findings indicate that Random Forest achieves the highest accuracy and F1-score for both valence and arousal, followed closely by AdaBoost and Decision Tree classifiers.
@@
Method Valence Arousal $uparrow Accuracy\uparrow F1-score\uparrow$ Nearest Neighbors 0.434 0.421 0.415 0.398 Linear SVM 0.474 0.419 0.488 0.417 RBF SVM 0.396 0.225 0.409 0.252 Decision Tree 0.484 0.461 0.525 0.486 Random Forest 0.526 0.497 0.547 0.476 AdaBoost 0.510 0.495 0.521 0.475 Naive Bayes 0.429 0.431 0.466 0.42 QDA 0.462 0.439 0.461 0.423
@@ IV-D Remote physiological signal extraction
Considering the results and importance of heart rate and respiration modalities in the emotion recognition of the previous subsection, we decide to extract these two signals remotely due to their potential link with emotional cues. To extract physiological signals from facial videos, we compare three state-of-the-art methods: two hand-crafted approaches, ICA [poh2010advancements] and POS [wang2016algorithmic], and one data-driven method, TDM [comas2022efficient]. The TDM model is a lightweight spatiotemporal network that effectively estimates rPPG spatiotemporal features by aggregating temporal derivative modules (TDM), emulating a Taylor series expansion. All these methods are chosen for their versatility, efficiency and suitability for real-time applications while maintaining competitive heart rate (HR) estimation performance. Table III presents the six-fold average performance of each method using common remote HR and respiration rate (RR) estimation metrics, including mean absolute error (MAE), root mean squared error (RMSE), and Pearson correlation ($\rho$). The results indicate that the hand-crafted models yield acceptable performance, with the POS method achieving an error of 2.01 BPM across the entire dataset. However, the data-driven TDM approach outperforms both, obtaining the lowest HR error with an MAE of 1.1 BPM, an RMSE of 4.51 BPM, and a correlation of 0.93. Additionally, although the original TDM model was designed exclusively for rPPG signal recovery, we enhanced it by, incorporating two final fully connected layers, yielding a multi-task regression framework that simultaneously estimates both rPPG and rRSP signals. The model was trained on the CAST-Phys dataset following the same six-fold cross-validation strategy described in Subsection IV-C. For remote respiration estimation, we achieved baseline results of 2.21 RPM MAE and 3.25 RPM RMSE.
Table IV provides detailed results across the six folds, showing that the TDM approach achieves consistently high accuracy in HR estimation, with MAE ranging from 0.72 to 1.71 BPM and strong correlations ($\rho>0.87$), particularly in Fold 4 ($\rho=0.98$). RR estimation exhibits greater variability, with MAE between 1.89 and 2.71 RPM and moderate correlations ($\rho=0.19–0.53$). Folds 1 and 4 yield the most stable performance for both HR and RR, while Folds 2, 3, and 5 show slightly degraded HR and RR correlations. Overall, these results demonstrate the robustness of the TDM method for HR extraction and its moderate sensitivity in RR estimation to inter-subject differences.
Given its superior performance and ability to extract both PPG and respiration signals, we employ the TDM model to obtain the remote physiological signals used for emotion recognition, as detailed in the next subsection. Figure 12 depicts an example of our remote physiological extraction during the high-arousal elicitation observed at the end of the video.
@@
TABLE IV: Remote heart rate and respiration estimation on the CAST-Phys using TDM approach for each fold.
[…]
  • MAE: Mean Absolute Error, RMSE: Root Mean Squared Error, $\rho$: Pearson Correlation
@@ IV-E Remote multi-modal emotion recognition
The selected features closely resemble those identified in Subsection IV-C. For the valence dimension, heart rate yields four significant features: pNN50, heart rate standard deviation, LF, and the LF/HF ratio, while respiration retains the same three meaningful features as before. Regarding arousal, heart rate includes four relevant features: LF, LF/HF ratio, TINN, and the Poincaré ratio between SD1 and SD2. Additionally, respiration now contributes two meaningful features: respiration rate mean and standard deviation.
@@
Figure 13 presents the accuracy performance of remote emotion recognition for each modality. For valence, heart rate exhibits a notable performance drop, achieving only 40.7% accuracy with Random Forest compared to the contact-based signal. In contrast, respiration maintains a similar performance, reaching 47.6% accuracy with Random Forest, closely matching the contact-based results. For arousal, the performance remains nearly unchanged between remote and contact-based modalities. Heart rate achieves 51.1% accuracy using Random Forest, slightly outperforming the 49.6% recorded in the contact-based setup. Similarly, respiration attains 46.3% accuracy with Random Forest, compared to 47.6% in the contact-based condition. These findings suggest that remote emotion recognition can achieve comparable results to contact-based methods, with the primary discrepancy observed in valence estimation using remote heart rate features.
@@
TABLE V: Remote multi-modal emotion recognition 6-Fold Average Performance
Method Valence Arousal $uparrow Accuracy\uparrow F1-score\uparrow$ Nearest Neighbors 0.404 0.395 0.370 0.351 Linear SVM 0.477 0.431 0.480 0.409 RBF SVM 0.396 0.225 0.409 0.244 Decision Tree 0.505 0.496 0.455 0.427 Random Forest 0.504 0.476 0.518 0.452 AdaBoost 0.485 0.470 0.49 0.443 Naive Bayes 0.455 0.453 0.454 0.408 QDA 0.426 0.424 0.445 0.413
@@
The drop in valence accuracy for remote heart rate can be attributed to the fewer features extracted by ANOVA, particularly the absence of important non-linear features that contribute to higher accuracy. In contrast, arousal performance benefits from the same two frequency-domain features used in the contact-based approach, along with two additional non-linear features, leading to a 1.5% improvement in accuracy. Hence, these results also highlight the critical role of feature selection in emotion recognition, as heart rate variability non-linear features appear to significantly impact both valence and arousal estimation. Exploring alternative feature selection methods or employing data-driven feature selection techniques could further enhance these baseline results.
@@
As shown in Subsection IV-C, the combination of the extracted modalities, in this case with facial expressions, rPPG and rRSP, reflects in Figure 13 and Table V a superior performance than each modality individually. From Table V, we appreciate that Random Forest, Decision Tree and Adaboost seem to be the most robust classifiers for the evaluated modalities like in contact-based estimation. Comparing the multi-modal emotion recognition results obtained from contact and fully remote modalities, can observe very similar performance with a slight decrease in remote modality. Figure 14 presents the confusion matrices for valence and arousal classification across different cases. For arousal (Figures 14c and 14d), we observe a common issue in both contact-based and remote approaches: the neutral state is not well detected, and arousal tends to be polarized between deactivated and activated emotions. This aligns with the findings discussed in previous subsections, where the presence of neutral states within activated and deactivated emotions makes accurate classification of this category more challenging. In contrast, for valence (Figures 14a and 14b), the neutral class achieves significantly better classification in both cases, with even stronger performance in remote emotion recognition. Although remote emotion recognition shows slightly lower accuracy for positive and negative states compared to contact-based methods, it appears to achieve a more balanced classification across the three arousal categories.
(a) Valence multi-modal contact-based
(b) Valence multi-modal remote-based
(c) Arousal multi-modal contact-based
(d) Arousal multi-modal remote-based
@@ V Discussion and future work
While this study presents the first approach to remote emotion estimation using three different modalities (rPPG, rRSP, facial expressions), EDA was not incorporated, as it remains largely unexplored in remote sensing. Recent studies [bhamborae2020towards, braun2024sympcam] have demonstrated the feasibility of recovering EDA signals through camera-based sensing. The integration of remote EDA into multi-modal emotion recognition frameworks could significantly enhance the estimation of sympathetic responses alongside modalities like rPPG or rRSP. In this sense, CAST-Phys can be a valuable dataset for such research lines, both in terms of providing training data and serving as a public benchmark.